OpenStudy (4everaddicted2anime):

Check my answer: Medal and Fan. An ice cream cone is 5" deep and 2" in top diameter. One scoop of vanilla ice cream, 2" in diameter, is placed on the cone. If the ice cream melts into the cone, will it overflow? Yes No I think it is yes, but I'm not sure

OpenStudy (4everaddicted2anime):

@FortyTheRapper @AnitaS.

OpenStudy (fortytherapper):

How did you approach this problem (Hint: Comparison of two volumes)

OpenStudy (4everaddicted2anime):

I need to compare volume of cone with volume of sphere V of cone=1/3 *pi*r^2*h V of sphere=4/3 *pi*r^3 V of cone= 1/3 * pi*1*5 = 5/3 pi V of sphere= 4/3 *pi*1 = 4/3 pi @FortyTheRapper the answer is no right?

OpenStudy (fortytherapper):

Perfect

OpenStudy (4everaddicted2anime):

Thank you!

OpenStudy (fortytherapper):

No problem